b3lla2006:

4y=6x-24 use the (y = mx + b) to solve

prettygirlgang:

wait 1 sec

prettygirlgang:

use this link https://www.mathway.com/popular-problems/Pre-Algebra/120544

supie:

We're graphing it or finding the slope?

b3lla2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie We're graphing it or finding the slope? \(\color{#0cbb34}{\text{End of Quote}}\) slope

b3lla2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @prettygirlgang use this link https://www.mathway.com/popular-problems/Pre-Algebra/120544 \(\color{#0cbb34}{\text{End of Quote}}\) thxs

prettygirlgang:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @b3lla2006 \(\color{#0cbb34}{\text{Originally Posted by}}\) @prettygirlgang use this link https://www.mathway.com/popular-problems/Pre-Algebra/120544 \(\color{#0cbb34}{\text{End of Quote}}\) thxs \(\color{#0cbb34}{\text{End of Quote}}\) np

jhonyy9:

4y = 6x -24 so to get this form of y = mx +b just divide both sides by 4 (4y)/4 = (6x)/4 -24/4 make the calcules hope helped understandably easy @Laylalyssa

jhonyy9:

@supie do you agree ?

supie:

Yes I think that's right.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Yes I think that's right. \(\color{#0cbb34}{\text{End of Quote}}\) ty

supie:

np